# How to Evaluate Distributed Coordination Systems? – A Survey and Analysis

### v2

@@ 2.2 Metrics
The number of regions of the deployment is a crucial factor for the wide area network systems due to the high network communication overhead. As mentioned in 2.1, in this survey only the real wide-area network environments or testbeds with compatible network latencies are considered separate regions. Although they are deployed over wide area networks, systems with a fixed size of deployments like WPaxos, WanKeeper, and ZooNet are not included for scalability evaluations for the number of regions due to the lack of sole measurements for the different number of regions. Omni-Paxos evaluates how it reacts to node failures and network partition with nodes in a fixed number of regions but does not discuss regional scalability. Among all the systems studied, only E-Paxos and Mencius purely analyze the system scalability for the number of regions. Mencius uses the fixed network latency by changing its regular topology from flat to star and having a central node to coordinate all network communications while E-Paxos uses EC2 placement groups for the deployment with varying network latencies.
Most systems studying scalability concentrate on workload scalability or horizontal system scalability, however, there are some exceptions. For instance, $M^{2}$ Paxos also studies the vertical scalability of the system by varying the CPU performance at the nodes. Similarly, PNUTS evaluates the effects of disk space on the average latency. Some systems evaluate scalability differently due to architectural variations. For instance, DistributedLog focused on scalability with respect to the number of streams being processed, since a stream is a basic unit of sharding in the system. However, because each physical node can only handle a limited amount of streams, scaling the workload to include more streams ultimately causes the underlying scheduling system to span additional worker nodes to handle the streams. DistributedLog evaluation addresses this by performing workload scalability benchmarks concerning a single proxy to study the scalability limitations of a single node while doing system scalability on a cluster that can span additional workers.
@@
Evaluating the consistency properties of a protocol or an algorithm is no trivial task. Part of the reason is understanding the consistency itself since many different consistency models exist. The ambiguity of consistency definitions also makes it difficult to compare the consistency guarantees provided by various systems without careful examination of the algorithms and protocols. For instance, many different systems claim strong consistency for their protocols, however, the actual guarantees provided may differ drastically, based on the transaction modes, assumptions about command ordering (total order, partial total order, etc), and assumptions about the client interaction with the system.
Although benchmarking serializability or linearizability is not simple, testing for these guarantees is feasible. Bizur and Multi-Paxos perform such tests to detect any inconsistencies in their strictly serialized write operations. Bizur uses Serialla, a testing tool for strict serializability in the Elastifile file system [6], which produces concurrent updates while checking the responses at all replicas. It detects the requests causing inconsistent execution orders at any replica and provides the descriptive log of operations. Similarly, Multi-Paxos uses runtime checking for any inconsistencies at any replica by periodically sending checksum requests to all replicas. Replicas calculate the checksum value of their fault-tolerant log and compare that with the master’s value. Jepsen [47] carries a test on ZooKeeper to confirm the linearizability variant maintained under network partition and leader failure by partitioning the site which has the leader and another replica and keeps sending the write requests during partition. Then, it recovers the region and checks the logs at each replica. Elle [48], built over Jepsen, efficiently checks for violations of ScalarDB’s strict serializability guarantees by constructing transaction dependency graphs and identifying critical dependency cycles.
In our study, we found very few systematic evaluations for consistency. However, some systems evaluate their consistency in terms of data staleness at different nodes or geographical regions. Such staleness-based evaluations aim to show that clients cannot read stale, or old, values of data, no matter which node is being used for reading the data. In this manner, ZooNet measures the performance degradation while disabling stale reads by synchronizing read requests first with the owner of the requested data object to make sure it serves the most up-to-date data.
@@
[…]
Regions
@@ C.S. ZooKeeper  [41] 5 1-2 Leader & Follower 1
The ability to tolerate failures and remain available is one of the properties of consensus algorithms and coordination systems built on top of such algorithms. The number of failures a system can mask often depends on the cluster size, and larger clusters can mask more failures. In Table IV, we summarize the vital parameters used by the systems when benchmarking for availability and fault tolerance. Most systems we have reviewed assumed no concurrent failures, however, ZooKeeper evaluation performed a benchmark with two follower nodes failing at almost the same time. The cluster size used for that evaluation of ZooKeeper allowed a maximum of two failures at the same time.
Spanner availability evaluation is different from the rest of the systems and protocols. Spanner uses Paxos-groups as a unit of data sharding and replication. In the availability evaluation, they created 1250 Paxos-groups on 25 nodes in 5 regions with each Paxos-group taking a single machine from each region. Then an entire region has been crashed, however, since Paxos protocol can mask up to 2 failures in the cluster of 5 nodes, all groups were able to continue operation. As a result, this benchmark is similar to the evaluation of availability on a single Paxos cluster of 5 nodes by crashing a node.
On the other hand, Hybrid Paxos and Tango approach availability evaluation differently and evaluate the effect of increased availability with respect to performance. Hybrid Paxos increases the number of replicas under the same load and measures the average latency to see the impact of having more replication in the system. Similarly, Tango uses a primary-backup scenario for the same view of the Tango object, serves all read requests from the backup replica and all writes from the primary replica, and measures the throughput accrodingly.
@@ 3 Benchmarking Requirements for Distributed Coordination Systems
The abundance of evaluations measuring the same aspects of systems behavior and the lack of a prevalent benchmarking suite suggests that existing benchmarking tools are either deficient in covering some aspects of evaluated metrics or are not universally adaptable. Identifying and addressing the requirements for a complete evaluation of distributed coordination systems will allow us to customize existing tools accordingly or to create a new benchmarking suite that can be used by a wide range of applications exposing similar interfaces to the user. In the remainder of this section, we will discuss benchmarking challenges and the general requirements of a benchmarking suite for distributed coordination systems.
@@ 3.1 Benchmarking Suite Flexibility and Sophistication
A flexible benchmark should be highly customizable throughout the range of parameters, but it also needs to allow for benchmarking various metrics of the system’s behavior. As such, a benchmarking suite for coordination systems needs to be able to evaluate all facets of the system’s behavior: performance, scalability, availability, and consistency. Some of these metrics are related, for instance, the benchmarking suite can evaluate the system’s scalability by measuring performance on a different system or workload. Some other metrics are orthogonal to each other, as is the case with performance and consistency.
@@
Read-to-write ratio is a fundamental property of workloads for coordination systems and applications relying on distributed synchronization. The amount of read and write operations a system is performing can vary greatly depending on the application. Many workloads are read-oriented, however, some tasks, such as logging, perform more updates.
Size of data pool and size of individual object control the overall size of the workload. Large object size can drastically increase the latency and reduce the throughput of the system, however, some systems optimize for larger objects while others work best with smaller data items. The size of the data pool is also used for generating variations for the percentage of conflicting commands along with the data access overlap ratio.
Data access overlap is important in the context of coordination services and coordinated applications. This parameter controls the likelihood of the same object being accessed by two or more different clients and hence it affects the command conflict ratio. In case of no overlap in data access, clients will never access objects belonging to other clients, while in a 100% overlap, all clients are equally likely to access any data. This measure is especially important for multi-leader coordination systems or applications that allow concurrent writes, where conflicting commands require a special resolution and a longer time to complete. This is also closely related to consistency models for those systems since for systems providing synchronous reads, all dependent requests need to be serialized. This is even more challenging for WAN deployments and has a significant impact on the latency evaluation.
@@
Scalability Benchmark. In addition to workload parameters, the scalability benchmarking suite needs to have the ability to tune the amount of work it pushes through the system under test. This is often achieved by increasing the number of concurrent clients interacting with the systems and/or increasing per-client command throughput. The size of the cluster, the number of servers, and the geo-distribution level are not the parameters of the benchmarking tool. However, the benchmarking tool should provide information about throughput/latency per node as well as the cumulative throughput/mean latency which enables comparing results from multiple runs of the benchmark with different configurations of the overall system architecture.
Availability Benchmark. Availability measurements require simulating system failures while measuring the performance of the system during them. Even though many types of failures are possible, all of the studied systems that did availability evaluation resorted to a crash-failure scenario. The ideal benchmarking suite needs to provide the ability to evaluate the system’s behavior not only under crash-failures, but under other types of malfunctions, such as network partitions, clock drifts, memory corruption, and unreliable links. For many of these failures, the benchmark needs to control the number of simultaneous failures.
Consistency Benchmark. Consistency evaluations significantly differ from the prior three benchmarks. Most importantly, consistency is not evaluated through performance observations. Existing attempts at consistency evaluations often focus on studying data-staleness. Staleness describes the amount of outdated data that can be read by the client. This is often good enough to show the eventual consistency guarantees of the system, but not nearly enough to test for all spectrums of possible consistency guarantees. This is especially important in the realm of consensus algorithms and coordination systems and applications using these algorithms since such systems aim to provide stronger levels of consistency than the eventual. In this matter, a consistency benchmark is required to evaluate the consistency guarantees provided by the system for different levels of linearizability and serializability under various combinations of client request ratios of read and write operations.
@@ 3.2 Benchmarking Suite for WAN Systems
Many of today’s distributed applications are deployed on scales that span multiple datacenters across the country, region, or even the globe. Such a scale introduces many challenges that are not observed in a single-datacenter setting. Large distances between components of the system drastically increase communication delay and thus systems’ latency. Cross-datacenter bandwidth may also be limited, driving performance degradation further.
Many WAN systems (such as [12, 65, 11]) also see the performance and scalability artifacts from the geographical placement of data centers. On a global scale, it is no longer possible to assume roughly uniform latencies between nodes located in different datacenters. Physical distances start to dictate the speed of communication between the regions, thus in a WAN system, the communication latency between regions can easily differ by the order of magnitude. These disproportional delays introduce penalties for some regions while giving benefits to others. In this matter, the distribution of the clients becomes also important since distributed clients may cause a variance in latency measurements. For instance, systems allowing their clients to communicate any replica at any region would differ in latency from the systems limiting client communication to local replicas.
@@ 3.4 Benchmarking Suite Ease of Adoption
The ease of use and adaptability of a benchmark is a big contributing factor to many systems deciding not to adopt any of the standard benchmarks for their evaluations. Straightforward integration for various systems developed with different programming languages and frameworks is critical for any benchmarking suite. It is also vital that the benchmarking suite should operate as a black box and should not require users to learn about the internals of the benchmarking suite. Similarly, the benchmarking suite should be configured for operating independently regardless of the evaluated system details such as programming languages used for the development. This is also an important factor for a fair comparison of the evaluated systems.
@@ 4 Existing Benchmarks and Their Deficiencies
TABLE V: Metrics provided by Benchmarking Tools
@@
TABLE VI: Parameters tunable by Benchmarking Tools
@@ 4.1 Performance
YCSB [23] has become a standard benchmarking tool for the evaluation of NoSQL frameworks since published in 2010. The main reason for its popularity is generality and extendability. YCSB is a general-purpose benchmarking tool and can be used for all NoSQL database systems. It is easy to extend to any additional database by implementing a simple CRUD (create, read, update, and delete) plus scan interface against the datastore under benchmarking. Although the YCSB workload does not model real applications, its generated synthetic workload is highly tunable in five dimensions: the number of clients, operation ratio, request distribution, key space, and throttled throughput. Even though it enables the distribution of the requests over key space to some level, it does not provide the requirements for evaluating access locality since it supports only single-client implementation. Additionally, despite the workload model supporting uniform, ziphian, latest, and multinominal distributions, it comes short of precise adjustments for data access patterns and it requires customization to set up with certain probabilities. Likewise, YCSB ignores creating conflicting commands. These parameters limit even the performance evaluations since systems have different consistency models and they have a significant impact on latency. For instance, a coordination system using a data ownership mechanism may need access migration for some data objects. It is simply not provided by any single-client benchmarking solution. YCSB only measures the performance metric in our category. In particular, it measures the latency of each operation and the overall throughput. Another limitation of YCSB is a lack of scalability for the WAN setting since YCSB only generates workload from one process. Our experience from benchmarking a geo-distributed system makes us realize that YCSB does not span over multiple datacenters and cannot generate workloads with some locality characteristics.
@@
YCSB+T [26] is an extension of YCSB that adds a new tier of transactional operations and validation. To generate a meaningful workload, YCSB+T defines the Closed Economy Workload which simulates bank account transactions with a fixed total amount. It supports transactional read, scan, update, delete, and readModifyWrite operations. The readModifyWrite operation includes reading two data objects and updating both. The validation phase of the benchmark tries to detect anomalies by comparing the total account balance before and after transactions. Given the database state, this validation cannot detect any dirty reads or lost updates that do not change the sum of all account balances. YCSB+T mimics the workload features of YCSB and it suffers from similar limitations. It is not well suited for WAN deployments and evaluating WAN systems as it does not support the distribution of the clients hence producing conflicting commands and managing access locality.
@@
Similarly, YCSB++ [70] extends YCSB to support multiple clients which can run on different machines. To manage synchronization and group membership of clients, it uses ZooKeeper as the coordination service and the notification mechanism. It occupies HBase [4] and Accumulo [3] as its table stores. To provide more realistic evaluations, it enables some other features of table stores such as table pre-splitting for fast ingest, server-side filtering, and bulk loading of the data. As another aspect of table stores, it enables evaluating the effect of applying access control on performance by using pre-configured access control lists (ACLs) to map credentials and operations for any schema. By using a monitoring tool, Otus [72], performance metrics of table stores and YCSB++ clients are collected in a central repository for fine-grained analysis. However, YCSB++ is tailored for table stores and more suited for evaluating big data applications. Moreover, the synchronization of the clients relies on the ZooKeeper service which performs poorly on the WAN scale. Although it supports the distribution of the clients, the distribution of the workload does not provide the ability to adjust data access overlap or access locality parameters.
@@
BG [15] is another benchmark that mainly focuses on performance under a specific real-world application workload. It models the workloads of social networking applications that have read operations like listing all friends or reading top posts for a user and write operations such as accepting friendship invitations. BG summarizes the performance evaluation in terms of Social Action Rating based on the customizable service level agreement (SLA). SLA is defined by four parameters, namely the percentage of requests to observe less than pre-specified response time, the response time, unpredictable data amount, and the unpredictable time limit for any data object. This simply models the measurements for latency and staleness. BG workload originally creates overlapping data access patterns since clients may request conflicting actions such as one client requesting acceptance of friendship invitations and another client requesting the rejection of the same invitation. However, BG uses a locking mechanism to avoid these conflicting commands. Similarly, although the clients can be distributed in BG, it only uses this for scalability measures and does not accommodate the need for access locality parameters.
@@
BenchFoundry [17] adopts a different approach to create workloads. Instead of providing predefined workloads, it allows the creation of the desired sequence of operations. In this way, it could be possible to generate custom workloads for the parameters listed in Table VI. However, it uses master-slave architecture before running the benchmarking clients which determines the order of the execution of the requests listed in each trace file. If we assume the jitter in clock synchronization is minimized, this would eliminate the conflicting commands even if clients can be configured to share the same key space.

### v3

@@ 2.2 Metrics
The number of regions of the deployment is a crucial factor for the wide area network systems due to the high network communication overhead. As mentioned in 2.1, in this survey, only the real wide-area network environments or testbeds with compatible network latencies are considered separate regions. Although they are deployed over wide area networks, systems with a fixed size of deployments like WPaxos, WanKeeper, and ZooNet are not included for scalability evaluations for the number of regions due to the lack of sole measurements for different numbers of regions. Omni-Paxos evaluates how it reacts to node failures and network partitions with nodes in a fixed number of regions, but does not discuss regional scalability. Among all the systems studied, only E-Paxos and Mencius purely analyze the system scalability for the number of regions. Mencius uses the fixed network latency by changing its regular topology from flat to star and having a central node to coordinate all network communications, while E-Paxos uses EC2 placement groups for the deployment with varying network latencies.
Most systems studying scalability concentrate on workload scalability or horizontal system scalability; however, there are some exceptions. For instance, $M^{2}$ Paxos also studies the vertical scalability of the system by varying the CPU performance at the nodes. Similarly, PNUTS evaluates the effects of disk space on the average latency. Some systems evaluate scalability differently due to architectural variations. For instance, DistributedLog focused on scalability with respect to the number of streams being processed, since a stream is a basic unit of sharding in the system. However, because each physical node can only handle a limited number of streams, scaling the workload to include more streams ultimately causes the underlying scheduling system to span additional worker nodes to handle the streams. DistributedLog evaluation addresses this by performing workload scalability benchmarks concerning a single proxy to study the scalability limitations of a single node while doing system scalability on a cluster that can span additional workers.
@@
Evaluating the consistency properties of a protocol or an algorithm is no trivial task. Part of the reason is understanding the consistency itself, since many different consistency models exist. The ambiguity of consistency definitions also makes it difficult to compare the consistency guarantees provided by various systems without careful examination of the algorithms and protocols. For instance, many different systems claim strong consistency for their protocols; however, the actual guarantees provided may differ drastically, based on the transaction modes, assumptions about command ordering (total order, partial total order, etc), and assumptions about the client interaction with the system.
Although benchmarking serializability or linearizability is not simple, testing for these guarantees is feasible. Bizur and Multi-Paxos perform such tests to detect any inconsistencies in their strictly serialized write operations. Bizur uses Serialla, a testing tool for strict serializability in the Elastifile file system [6], which produces concurrent updates while checking the responses at all replicas. It detects the requests causing inconsistent execution orders at any replica and provides a descriptive log of operations. Similarly, Multi-Paxos uses runtime checking for any inconsistencies at any replica by periodically sending checksum requests to all replicas. Replicas calculate the checksum value of their fault-tolerant log and compare that with the master’s value. Jepsen [49] carries out a test on ZooKeeper to confirm the linearizability variant maintained under network partition and leader failure by partitioning the site that has the leader and another replica and keeps sending the write requests during partition. Then, it recovers the region and checks the logs at each replica. Elle [50], built over Jepsen, efficiently checks for violations of ScalarDB’s strict serializability guarantees by constructing transaction dependency graphs and identifying critical dependency cycles.
In our study, we found very few systematic evaluations for consistency. However, some systems evaluate their consistency in terms of data staleness at different nodes or geographical regions. Such staleness-based evaluations aim to show that clients cannot read stale or old values of data, no matter which node is being used for reading the data. In this manner, ZooNet measures the performance degradation while disabling stale reads by synchronizing read requests first with the owner of the requested data object to make sure it serves the most up-to-date data.
@@
System # of Servers # of Failures        Failed Node(s)Type     # ofRegions
@@ C.S. ZooKeeper  [43] 5 1-2 Leader & Follower 1
The ability to tolerate failures and remain available is one of the properties of consensus algorithms and coordination systems built on top of such algorithms. The number of failures a system can mask often depends on the cluster size, and larger clusters can mask more failures. In Table IV, we summarize the vital parameters used by the systems when benchmarking for availability and fault tolerance. Most systems we have reviewed assumed no concurrent failures; however, ZooKeeper evaluation performed a benchmark with two follower nodes failing at almost the same time. The cluster size used for that evaluation of ZooKeeper allowed a maximum of two failures at the same time.
Spanner availability evaluation is different from the rest of the systems and protocols. Spanner uses Paxos groups as a unit of data sharding and replication. In the availability evaluation, they created 1250 Paxos groups on 25 nodes in 5 regions, with each Paxos group taking a single machine from each region. Then an entire region has been crashed; however, since the Paxos protocol can mask up to 2 failures in a cluster of 5 nodes, all groups were able to continue operation. As a result, this benchmark is similar to the evaluation of availability on a single Paxos cluster of 5 nodes by crashing a node.
On the other hand, Hybrid Paxos and Tango approach availability evaluation differently and evaluate the effect of increased availability with respect to performance. Hybrid Paxos increases the number of replicas under the same load and measures the average latency to see the impact of having more replication in the system. Similarly, Tango uses a primary-backup scenario for the same view of the Tango object, serves all read requests from the backup replica and all writes from the primary replica, and measures the throughput accordingly.
@@ 3 Benchmarking Requirements for Distributed Coordination Systems
The abundance of evaluations measuring the same aspects of system behavior and the lack of a prevalent benchmarking suite suggest that existing benchmarking tools are either deficient in covering some aspects of evaluated metrics or are not universally adaptable. Identifying and addressing the requirements for a complete evaluation of distributed coordination systems will allow us to customize existing tools accordingly or to create a new benchmarking suite that can be used by a wide range of applications exposing similar interfaces to the user. In the remainder of this section, we will discuss benchmarking challenges and the general requirements of a benchmarking suite for distributed coordination systems.
@@ 3.1 Benchmarking Suite Flexibility and Sophistication
A flexible benchmark should be highly customizable throughout the range of parameters, but it also needs to allow for benchmarking various metrics of the system’s behavior. As such, a benchmarking suite for coordination systems needs to be able to evaluate all facets of the system’s behavior: performance, scalability, availability, and consistency. Some of these metrics are related; for instance, the benchmarking suite can evaluate the system’s scalability by measuring performance on a different system or workload. Some other metrics are orthogonal to each other, as is the case with performance and consistency.
@@
Read-to-write ratio is a fundamental property of workloads for coordination systems and applications relying on distributed synchronization. The number of read and write operations a system is performing can vary greatly depending on the application. Many workloads are read-oriented; however, some tasks, such as logging, perform more updates.
Size of data pool and size of individual object control the overall size of the workload. Large object size can drastically increase the latency and reduce the throughput of the system; however, some systems optimize for larger objects while others work best with smaller data items. The size of the data pool is also used for generating variations for the percentage of conflicting commands, along with the data access overlap ratio.
Data access overlap is important in the context of coordination services and coordinated applications. This parameter controls the likelihood of the same object being accessed by two or more different clients and hence it affects the command conflict ratio. In case of no overlap in data access, clients will never access objects belonging to other clients, while in a 100% overlap, all clients are equally likely to access any data. This measure is especially important for multi-leader coordination systems or applications that allow concurrent writes, where conflicting commands require a special resolution and a longer time to complete. This is also closely related to consistency models for those systems, since for systems providing synchronous reads, all dependent requests need to be serialized. This is even more challenging for WAN deployments and has a significant impact on the latency evaluation.
@@
Scalability Benchmark. In addition to workload parameters, the scalability benchmarking suite needs to have the ability to tune the amount of work it pushes through the system under test. This is often achieved by increasing the number of concurrent clients interacting with the systems and/or increasing per-client command throughput. The size of the cluster, the number of servers, and the geo-distribution level are not the parameters of the benchmarking tool. However, the benchmarking tool should provide information about throughput/latency per node as well as the cumulative throughput/mean latency, which enables comparing results from multiple runs of the benchmark with different configurations of the overall system architecture.
Availability Benchmark. Availability measurements require simulating system failures while measuring the performance of the system during them. Even though many types of failures are possible, all of the studied systems that underwent availability evaluation resorted to a crash-failure scenario. The ideal benchmarking suite needs to provide the ability to evaluate the system’s behavior not only under crash-failures, but under other types of malfunctions, such as network partitions, clock drifts, memory corruption, and unreliable links. For many of these failures, the benchmark needs to control the number of simultaneous failures.
Consistency Benchmark. Consistency evaluations significantly differ from the prior three benchmarks. Most importantly, consistency is not evaluated through performance observations. Existing attempts at consistency evaluations often focus on studying data staleness. Staleness describes the amount of outdated data that can be read by the client. This is often good enough to show the eventual consistency guarantees of the system, but not nearly enough to test for all spectrums of possible consistency guarantees. This is especially important in the realm of consensus algorithms and coordination systems and applications using these algorithms, since such systems aim to provide stronger levels of consistency than eventual. In this matter, a consistency benchmark is required to evaluate the consistency guarantees provided by the system for different levels of linearizability and serializability under various combinations of client request ratios of read and write operations.
@@ 3.2 Benchmarking Suite for WAN Systems
Many of today’s distributed applications are deployed on scales that span multiple datacenters across the country, region, or even the globe. Such a scale introduces many challenges that are not observed in a single-datacenter setting. Large distances between components of the system drastically increase communication delay and thus the system’s latency. Cross-datacenter bandwidth may also be limited, driving performance degradation further.
Many WAN systems (such as [12, 67, 11]) also see the performance and scalability artifacts from the geographical placement of data centers. On a global scale, it is no longer possible to assume roughly uniform latencies between nodes located in different datacenters. Physical distances start to dictate the speed of communication between the regions; thus, in a WAN system, the communication latency between regions can easily differ by the order of magnitude. These disproportional delays introduce penalties for some regions while giving benefits to others. In this matter, the distribution of the clients also becomes important since distributed clients may cause a variance in latency measurements. For instance, systems allowing their clients to communicate with any replica at any region would differ in latency from the systems limiting client communication to local replicas.
@@ 3.4 Benchmarking Suite Ease of Adoption
The ease of use and adaptability of a benchmark is a big contributing factor to many systems deciding not to adopt any of the standard benchmarks for their evaluations. Straightforward integration for various systems developed with different programming languages and frameworks is critical for any benchmarking suite. It is also vital that the benchmarking suite operates as a black box and does not require users to learn about the internals of the benchmarking suite. Similarly, the benchmarking suite should be configured to operate independently regardless of the evaluated system details, such as the programming languages used for the development. This is also an important factor for a fair comparison of the evaluated systems.
@@ 4 Existing Benchmarks and Their Deficiencies
TABLE V: Metrics provided by benchmarking tools.
@@
TABLE VI: Parameters tunable by benchmarking tools.
@@ 4.1 Performance
YCSB [24] has become a standard benchmarking tool for the evaluation of NoSQL frameworks since its publication in 2010. The main reason for its popularity is generality and extendability. YCSB is a general-purpose benchmarking tool and can be used for all NoSQL database systems. It is easy to extend to any additional database by implementing a simple CRUD (create, read, update, and delete) plus scan interface against the datastore under benchmarking. Although the YCSB workload does not model real applications, its generated synthetic workload is highly tunable in five dimensions: the number of clients, operation ratio, request distribution, key space, and throttled throughput. Even though it enables the distribution of the requests over the key space to some level, it does not provide the requirements for evaluating access locality since it supports only a single-client implementation. Additionally, despite the workload model supporting uniform, ziphian, latest, and multinomial distributions, it comes short of precise adjustments for data access patterns, and it requires customization to set up with certain probabilities. Likewise, YCSB ignores creating conflicting commands. These parameters limit even the performance evaluations since systems have different consistency models, and they have a significant impact on latency. For instance, a coordination system using a data ownership mechanism may need access migration for some data objects. It is simply not provided by any single-client benchmarking solution. YCSB only measures the performance metric in our category. In particular, it measures the latency of each operation and the overall throughput. Another limitation of YCSB is a lack of scalability for the WAN setting since YCSB only generates workload from one process. Our experience from benchmarking a geo-distributed system makes us realize that YCSB does not span multiple datacenters and cannot generate workloads with some locality characteristics.
@@
YCSB+T [28] is an extension of YCSB that adds a new tier of transactional operations and validation. To generate a meaningful workload, YCSB+T defines the Closed Economy Workload, which simulates bank account transactions with a fixed total amount. It supports transactional read, scan, update, delete, and readModifyWrite operations. The readModifyWrite operation includes reading two data objects and updating both. The validation phase of the benchmark tries to detect anomalies by comparing the total account balance before and after transactions. Given the database state, this validation cannot detect any dirty reads or lost updates that do not change the sum of all account balances. YCSB+T mimics the workload features of YCSB, and it suffers from similar limitations. It is not well-suited for WAN deployments and evaluating WAN systems, as it does not support the distribution of the clients, hence producing conflicting commands and managing access locality.
@@
Similarly, YCSB++ [72] extends YCSB to support multiple clients that can run on different machines. To manage synchronization and group membership of clients, it uses ZooKeeper as the coordination service and the notification mechanism. It occupies HBase [4] and Accumulo [3] as its table stores. To provide more realistic evaluations, it enables some other features of table stores, such as table pre-splitting for fast ingest, server-side filtering, and bulk loading of the data. As another aspect of table stores, it enables evaluating the effect of applying access control on performance by using pre-configured access control lists (ACLs) to map credentials and operations for any schema. By using a monitoring tool, Otus [74], performance metrics of table stores and YCSB++ clients are collected in a central repository for fine-grained analysis. However, YCSB++ is tailored for table stores and is more suited for evaluating big data applications. Moreover, the synchronization of the clients relies on the ZooKeeper service, which performs poorly on the WAN scale. Although it supports the distribution of the clients, the distribution of the workload does not provide the ability to adjust data access overlap or access locality parameters.
@@
BG [16] is another benchmark that mainly focuses on performance under a specific real-world application workload. It models the workloads of social networking applications that have read operations like listing all friends or reading top posts for a user, and write operations such as accepting friendship invitations. BG summarizes the performance evaluation in terms of Social Action Rating based on the customizable service level agreement (SLA). SLA is defined by four parameters, namely the percentage of requests to observe less than pre-specified response time, the response time, the unpredictable data amount, and the unpredictable time limit for any data object. This simply models the measurements for latency and staleness. BG workload originally creates overlapping data access patterns since clients may request conflicting actions, such as one client requesting acceptance of friendship invitations and another client requesting the rejection of the same invitation. However, BG uses a locking mechanism to avoid these conflicting commands. Similarly, although the clients can be distributed in BG, it only uses this for scalability measures and does not accommodate the need for access locality parameters.
@@
BenchFoundry [18] adopts a different approach to create workloads. Instead of providing predefined workloads, it allows the creation of the desired sequence of operations. In this way, it could be possible to generate custom workloads for the parameters listed in Table VI. However, it uses master-slave architecture before running the benchmarking clients, which determines the order of the execution of the requests listed in each trace file. If we assume the jitter in clock synchronization is minimized, this would eliminate the conflicting commands even if clients can be configured to share the same key space.
HiBench [42] aims to provide a more comprehensive and realistic evaluation of Hadoop with various pre-defined workloads such as web searches, machine learning tasks, and HDFS operations. The workloads offer different object sizes and data access patterns, but are not effective in evaluating conflicting operations by nature of the Hadoop workflow, and apply only to Hadoop-like systems. BigDataBench [83] expands on this with more workloads and targeted systems, but does not allow users to tune conflicts.